OpenStudy (anonymous):

write an equation for the line that is perpendicular to the given line and that passes through the given point. y= -1/3x + 2 ; (4,2) the answer to the equation is : y=3x-10 but i don't understand how it got "-10" my original answer was y=3x-2 . please help!

OpenStudy (turingtest):

Sorry, transposed the x and y new slope: m=3 y-2=3(x-4) y=3x-10 I'm tired :(

OpenStudy (anonymous):

ok thank you!